A faucet is leaking 60 drops per minute. Determine amount of water wasted in 30 days.
A drop is approx. 0.05 milliliters.
Not sure how to solve.

Found 3 solutions by josgarithmetic, mananth, ikleyn:
Answer by josgarithmetic(39617)

Treat the rate as or with unit conversions.

60 drops per minute; how much in 30 days?



Simplify and compute that.



130 liters in thirty days

Answer by mananth(16946)

A faucet is leaking 60 drops per minute. Determine amount of water wasted in 30 days.
A drop is approx. 0.05 milliliters.
1 day = 24*60 minutes
In 1 minute 60 drops
in (24*60) minutes number of drops = 24*60*60 drops
A drop is approx. 0.05 milliliters.
24*60*60 drops = 24*60*60*0.05 mL
= 4320 mL wasted in 30 days


Answer by ikleyn(52781)

.
A faucet is leaking 60 drops per minute. Determine amount of water wasted in 30 days.
A drop is approx. 0.05 milliliters.
Not sure how to solve.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        Calculations in the post by @mananth are  INCORRECT  (the factor of  30  days is missed there).
        So,  I came to bring a correct solution.


60 drops per minute provide 60*0.05 = 3 mL.

The day has 24*60 = 1440 minutes.


So, the waste per day is 3*1440 = 4320 mL.


It should be multiplied by 30 days, giving total waste of water  30*4320 = 129600 mL,

or 129.6 liters of water.


ANSWER.  129.6 liters of water wasted in 30 days.

Solved.

These calculations can be done in one line
(60  at the second appearance is the number of minutes in one hour;  24  is the number of hours in a day;
the rest is clear without explanations)

        waste volume = 0.05 * 60 * 60 * 24 * 30 = 129,600 mL = 129.6 Liters.